# Win Ratio with Multiple Thresholds for Composite Endpoints

### v4

@@ 3.3 Influence of the correlation
In this subsection, we present additional insights into the observed difference in power between simulation scenarios S3 and S4. The win ratios (full spelling for the ratio measures themselves, and WR for the method) in simulation scenarios S3 and S4 with 750-day follow-up time are decomposed to stage-level summaries and averaged across 2000 replicates in Table 3. Results from other follow-up times are included in Appendix C. The decomposition procedure presents the wins, losses, and ties obtained by the treatment group participants in each stage, where the ratio of the number of such wins to the number of such losses corresponds to the stage-level win ratio. For example, for WR under simulation scenario S3 in Table 3, the averaged wins (losses) obtained in the survival time layer is 37.29% (27.56%). The 35.15% pairs with ties on the survival time layer then get their time-to-hospitalization endpoint compared, resulting in 14.93% wins and 17.07% losses. The remaining 3.15% pairwise comparisons receive ties as their final conclusions. In such a decomposition matrix, the percentages are all calculated based on the total number of pairwise comparisons between two contrasting participants. Therefore, the overall win ratio can be obtained by dividing the sum of stage-level wins by the sum of stage-level losses. In Table 3, the bold values in the stage-level win ratio column are those obtained by the time-to-hospitalization endpoint. In simulation scenario S3, these values are less than one, as opposed to very close to one in simulation scenario S4. Importantly, the difference between around 0.87 and 1.00 in terms of the stage-level win ratios averaged across 2000 replicates is non-negligible. In simulation scenario S3, the marginal correlation between the survival time and time-to-hospitalization endpoints influences the chances to actually compare the time-to-hospitalization endpoint (conditioning on receiving ties in the survival time layer). As a result, although there is no treatment effect on the time-to-hospitalization endpoint marginally, the spurious, negative “treatment effects” are observed based on the pairwise comparisons between participants with no difference on the survival time layer. Such spurious negative “treatment effects” are the core reason for the reduced power in detecting the true positive treatment effects with both WR and WR-AT. These results underscore the importance of carefully considering the inclusion of additional endpoints in a hierarchical testing structure, as inter-endpoint correlations can affect treatment effect detection in unanticipated ways (Verbeeck et al., 2019). An endpoint that appears neutral with respect to power may, after conditioning on higher-layer outcomes, exhibit spurious negative or positive “treatment effects”, potentially affecting statistical inference. Additionally, Table 3 shows that the final stages of WR and WR-AT leave the same proportion of ties (3.15% in simulation scenario S3 and 1.34% in S4, respectively). This confirms that setting zero in the last two stages of WR-AT prevents introducing additional ties and ensures full utilization of information, as in the standard WR.
@@
Table 3: Stage-level decomposition of win ratios in simulation scenarios S3 and S4 with 750-day follow-up time averaged across 2000 replicates. $\mathscr{D}_{ij}(d)$ and $\mathscr{T}_{ij1}(t_{1})$ represent comparison functions for survival time and time-to-hospitalization endpoints with thresholds $d$ and $t_{1}$, respectively. Similar for $\mathscr{D}_{ij}(0)$ and $\mathscr{T}_{ij1}(0)$. The stage-level win ratios with the bolded values correspond to those relevant to the time-to-hospitalization endpoint. Both simulation scenarios S3&S4 feature modest treatment effects on the survival time endpoint only, where simulation scenario S3 assumes a correlation between endpoints and simulation scenario S4 assumes no such correlation.
Scenario Method Stage Win (%) Tie (%) Loss (%) Stage-level win ratio
S3 WR $\mathscr{D}_{ij}(0)$ 37.29 35.15 27.56 1.35
$\mathscr{T}_{ij1}(0)$ 14.93 3.15 17.07 0.87
WR-AT $\mathscr{D}_{ij}(d)$ 30.27 48.01 21.71 1.39
$\mathscr{T}_{ij1}(t_{1})$ 16.00 13.28 18.73 0.85
$\mathscr{D}_{ij}(0)$ 2.48 8.76 2.04 1.21
$\mathscr{T}_{ij1}(0)$ 2.77 3.15 2.85 0.97
S4 WR $\mathscr{D}_{ij}(0)$ 37.22 35.18 27.60 1.35
$\mathscr{T}_{ij1}(0)$ 16.90 1.34 16.94 1.00
WR-AT $\mathscr{D}_{ij}(d)$ 30.20 48.04 21.76 1.39
$\mathscr{T}_{ij1}(t_{1})$ 18.31 11.39 18.34 1.00
$\mathscr{D}_{ij}(0)$ 2.39 6.86 2.14 1.12
$\mathscr{T}_{ij1}(0)$ 2.76 1.34 2.76 1.00
@@ 3.4 Influence of the caliper
In this subsection, we demonstrate the influence of caliper choices on adaptive thresholds. In addition to the default caliper ($c=20\%$), we consider $c=10\%,40\%$, and the combined calipers. The combined caliper is an eight-stage WR-MT that progressively employs $c=40\%,20\%,10\%$ calipers: $\text{WR-MT}\left(q_{D}^{40\%},q_{1}^{40\%},q_{D}^{20\%},q_{1}^{20\%},q_{D}^{10\%},q_{1}^{10\%},0,0\right).$ The results are presented in Table 4. In this simulation, a larger caliper is more effective at bypassing random differences, while a smaller caliper is more sensitive to detecting treatment effects on the survival time layer. The combined caliper is slightly enhanced by the additional stages used to bypass random differences. When increasing the caliper from 10% to 20% and 40%, the performance gains in simulation scenarios S1, S2, S5, and S6, where treatment effects primarily occur on the time-to-hospitalization layer, generally outweigh the performance losses observed in the remaining scenarios where effects are concentrated on the survival time layer. This is expected since the adaptive thresholds work to keep detecting true signals while bypassing random differences. The combined caliper approach achieves the highest power under simulation scenarios S1, S2, S5, and S6, which is likely a result of its enhanced comparison structure that is more effective at bypassing the random differences. In general, WR-AT remains robust to the choice of caliper values as no rapid change in the performance of WR-AT with different calipers is observed under most simulation scenarios.
Table 4: Empirical power (%) obtained by WR-AT with different calipers $c$ in simulation scenarios S1-S8. Notation: FU represents the schedule follow-up length; Caliper-10% refers to WR-AT analysis with caliper $c=10\%$, similar for 20% and 40%; Combined Caliper refers to an eight-stage WR-AT that progressively employs $c=40\%,20\%,10\%$ calipers.
@@ 3.5 Influence of the weight
In this subsection, we show the influence of weight choices on adaptive thresholds. Alternative to the default weight choice $w=1$, we also examine $w=0.1,0.3,0.5$. In the resulting four candidate methods, WR-AT(1), WR-AT(0.5), WR-AT(0.3), WR-AT(0.1), a smaller $w$ places greater emphasis on the survival time endpoint. According to the results presented in Table 5, in simulation scenarios S1, S2, S5, and S6, where the treatment effects are primarily on the time-to-hospitalization layer, smaller $w$ leads to lower power. This is not surprising since emphasizing the higher-layer endpoint naturally reduces the chance in detecting the treatment effect on the lower layer. In simulation scenarios S3, S4, S7, and S8, where the treatment effects are primarily on the survival time layer, relatively more comparable power is achieved by different choices of $w$. These indicate that WR-AT is generally robust to the choice of $w$ in terms of maintaining the capacity to detect treatment effects on the survival time layer.
Table 5: Empirical power (%) obtained by WR-AT with different weights $w$ in simulation scenarios S1-S8. Notation: FU represents the schedule follow-up length; WR-AT(0.1) refers to WR-AT analysis with weight $w=0.1$, similar for WR-AT(0.3), WR-AT(0.5), and WR-AT(1.0).
Scenario FU WR-AT(0.1) WR-AT(0.3) WR-AT(0.5) WR-AT(1.0)
@@ 4 Case Study: Digitalis Investigation Group (DIG) clinical trial
We apply WR-AT to analyze the Digitalis Investigation Group (DIG) trial (ClinicalTrials.gov identifier: NCT00000476). This clinical trial investigated the effect of digoxin on patients with heart failure. Patients who had heart failure and a left ventricular ejection fraction of 0.45 or less were eligible for the primary randomization. This study was neutral for its primary outcome of all-cause mortality but evaluated both death and hospitalization events. When these two outcomes were analyzed separately, treatment effects appeared mainly on the hospitalization outcomes. Alternatively, we hereby employ WR-AT and standard WR to reanalyze this dataset. In this case study, both death and hospitalization outcomes are analyzed in the time-to-event form with the right censoring information taken into account (i.e., survival time and time to hospitalization). In developing the endpoint hierarchy, time to (all-cause) death is included as the higher layer, and time to (all-cause) hospitalization is included as the lower layer in both WR-AT and WR. To bring new insight into this dataset, we focus on key subgroups from the trial population and use stratification for better adjustment in this analysis. To be specific, we focus on the subgroup population with New York Heart Association (NYHA) functional class III or IV and stratify by ejection fraction (<0.25<0.25 or 0.250.450.25-0.45), cause of heart failure (ischemic or nonischemic), and age (<70<70 or $\geq 70$). The number of patients within each stratum is presented in Appendix D. For WR-AT, the adaptive thresholds with default weight and caliper ($w=1,c=20\%$) are adopted. In both approaches, stratification follows Dong et al. (2018) for the calculation of win ratios and Finkelstein and Schoenfeld (1999) for hypothesis testing. The results are shown in Table 6.
Table 6: Hypothesis testing results of WR-AT and WR for the case study
Win ratio P-value
WR-AT 1.161 0.033
WR 1.134 0.055
@@
Both methods result in similar estimates of the win ratio, 1.161 with WR-AT and 1.134 with WR. However, WR-AT yields a smaller p-value and concludes a significant treatment effect. It is noticeable that, the treatment effect is weak on the survival time endpoint but relatively stronger on the time-to-hospitalization endpoint; therefore WR may be less powerful as it assigns the strict priority given to the survival time. In contrast, WR-AT demonstrates a higher capacity in detecting the treatment effect under such cases. Finally, we report p-values for the hypothesis tests and interpret statistical significance at the 0.05 level only for the purposes of demonstration. Nevertheless, we recognize the caveat that, in practice, study conclusions may not depend solely on whether a p-value is greater than or less than a single cutoff point like 0.05.
@@ 5 Discussion
In this study, we develop a hypothesis testing approach for composite time-to-event endpoints with a hierarchical structure. Modifying the standard WR, the proposed WR-MT adds layers with non-zero thresholds to pairwise comparisons. WR-MT gains power over the standard WR by bypassing random differences on the higher layer to allow meaningful contributions from the lower layers. Weighted adaptive thresholds are also developed to reduce the potential challenge in determining the non-zero thresholds. With simulated datasets, we compare the performance of WR-AT and WR while addressing the influence of the follow-up time, treatment effect sizes, and correlations between the two endpoints. The numerical results show that WR-AT outperforms WR when the effect is primarily on the lower layer and generally maintains comparable power when the effect is primarily on the higher layer.
We emphasize that, whether using WR or WR-MT (WR-AT), clinical relevance remains central in selecting and prioritizing endpoints in the outcome specification stage. In general, WR-MT still assumes that composite endpoints and their hierarchy are defined a priori based on clinical judgment. That is, the WR-MT framework does not completely overturn this ordering; rather, it introduces a more flexible comparison mechanism that starts from the highest-priority endpoint but then permits evaluation of lower-tier endpoints only when differences on the top layer are within a non-informative range. Conceptually, WR-MT occupies a middle ground between strict hierarchical enforcement (standard WR) and complete re-ranking of endpoints, and introduces a formal approach to balances clinical relevance with statistical efficiency. In practice, if strong clinical priors exist for fixed thresholds, they should be incorporated within WR-MT. In the absence of such priors, WR-AT offers a pragmatic mechanism to adaptively define these thresholds. We have not argued the use of WR-MT and WR-AT to replace standard WR in all settings. In fact, when clinical relevance demands strict hierarchical testing, standard WR will still be preferred as primary analysis but WR-MT and WR-AT may contribute useful information in secondary analysis. In any case, large discrepancies between standard WR and WR-MT (WR-AT) results should prompt reflection on what aspect of treatment benefit is most meaningful given the context of the specific clinical trial.
@@
In WR-AT, there are two tuning parameters: weight $w$ and caliper $c$. As a general rule of thumb, we recommend using $w=1,c=20\%$ as the default when clinical priors are not sufficiently strong. We have investigated several different choices of $c$ in the simulations to provide additional insights into their working mechanisms. In addition to different choices of $c$, the combined caliper approach is a strong alternative (e.g., $\text{WR-MT}\left(q_{D}^{40\%},q_{1}^{40\%},q_{D}^{20\%},q_{1}^{20\%},q_{D}^{10\%},q_{1}^{10\%},0,0\right)$ as employed in Section 3.4). The performance of the combined caliper approach tends to rely mainly on the largest caliper because the priority is given to this caliper by the specification of the hierarchy. It is also worth noting that extra attentions are given to the lower layer (i.e., time-to-hospitalization endpoint in our example) in this approach since each one of the added calipers will provide additional opportunities for the lower layer to contribute information to the pairwise comparison. In this combined caliper approach, the additional computation time brought by the added stages may need to be considered, particularly when the study sample size is large or repeated calculations are required (e.g., resampling-based inference). On the other hand, when $c$ is fixed, the weight $w\in[0,1]$ increases the adaptive threshold of the lower layer to address the higher layer endpoint. Since the reciprocal function in the $q_{1}^{c}/w$ structure is not on a linear scale, care is needed to use a relatively small $w$. Apart from adopting the reciprocal function, one may also consider the two-caliper structure ($c_{1},c_{2}$). This architecture employs a larger quantile to increase the lower layer threshold. When a small $w$ is desired, one may employ this approach and adjust $c_{2}$ to avoid the potential trivial stage when $w$ is specified close to zero. Despite the flexibility of the proposed adaptive threshold system, the structure ($w,c$) or ($c_{1},c_{2}$) should be pre-specified. Finally, when domain knowledge allows for precisely specifying all thresholds, directly implementing the WR-MT structure could potentially enhance the clinical interpretability of each stage. In general, WR-AT serves to simplify the process for selecting tuning parameters based on general rule of thumbs, when clinical information is insufficient to determine the thresholds. In trials where there is a preference for strict pre-specification, WA-AT may be recommended for secondary analysis.
Although we have primarily focused on the win ratio measure, we acknowledge the ongoing debate surrounding the statistical interpretation of the win ratio measure, particularly its handling of tied pairs and the implications for estimation and inference. Recent critiques, including Verbeeck et al. (2023), have highlighted the limitations of the win ratio measure in addressing ties, and have recommended difference-based measures, such as net treatment benefit or probability index, as more interpretable alternatives. Additionally, the win odds measure has been suggested as a viable alternative that overcomes some of the limitation of win ratio measure (Gasparyan et al., 2021; Brunner et al., 2021). In general, the proposed approach can be extended beyond the win ratio measure. That is, the key idea of establishing comparison functions (e.g., $\mathscr{D}_{ij}(d),\mathscr{T}_{ij1}(t_{1})$) by applying fixed or adaptive thresholds successively and alternating across endpoints to release the strict hierarchy can be naturally adapted to create alternative win statistics, such as win probability (Gasparyan et al., 2021) and net treatment benefit (Buyse, 2010) with multiple thresholds. Finally, the win odds with multiple thresholds can be defined by modifying equation (1) through adding $1/2$ of the tied pairwise comparisons to both the numerator and denominator. However, a systematic exploration of the operating characteristics of these additional win statistics with multiple, possibly adaptive, thresholds is beyond the scope of this paper and will be left for future work.
There are other potential extensions that we may address in future studies. First, similar to the standard WR, the estimated win ratios in WR-MT are affected by right censoring of time-to-event endpoints (Oakes, 2016). The existing methods to address right censoring such as inverse probability of censoring weighting (Dong et al., 2020, 2021) may be extended to WR-MT, and are worth investigation in future research. Second, in addition to the time-to-event endpoints, it is possible to include other types of outcome measures (e.g., continuous outcomes) in WR-MT. Finally, by employing the comparison functions defined in WR-MT, the multiple thresholds method can be extended to the existing regression frameworks developed for win functions (Mao and Wang, 2021; Wang and Mao, 2022; Song et al., 2023). Such extensions may help explore associational effects in clinical trials and observational studies and represent fruitful directions for future studies.

### v5

@@ 3.3 Influence of the correlation
In this subsection, we present additional insights into the observed difference in power between simulation scenarios S3 and S4 by examining how each endpoint or stage contributes through decomposition matrices. The decomposition procedure presents the wins, losses, and ties for the treatment group participants in each stage. In such a decomposition matrix, the percentages are all calculated based on the total number of pairwise comparisons between two contrasting participants. Then the calculation of three GPC measures at the stage level, NB, WR, and WO, follows. Results with 750-day follow-up time averaged across 2000 replicates are shown in Table 3. Results from other follow-up times are included in Supporting Information S3.
In our example, the correlation between endpoints is more likely to influence the overall testing through the time-to-hospitalization endpoint, since the survival time endpoint is prioritized and therefore less likely to be affected by the correlation at the first stage. To evaluate the contribution of the time-to-hospitalization endpoint in the FS test, we naturally consider the only stage $\mathscr{T}_{ij1}(0)$, which directly reflects this component’s effect. For the FS-AT test, between the stages $\mathscr{T}_{ij1}(t_{1})$ and $\mathscr{T}_{ij1}(0)$, we give more attention to $\mathscr{T}_{ij1}(t_{1})$ stage, where more ties are resolved by the time-to-hospitalization endpoint, leading to a greater influence on the overall testing outcome. In simulation scenario S3, NB is negative and both WO and WR are less than one on these stages, suggesting potential negative stage-level treatment effects, whereas in simulation scenario S4, they are very close to zero or one, indicating no treatment effects. Importantly, in the FS test, the difference between approximately -6.10% in simulation scenario S3 and -0.10% in simulation scenario S4 for stage-level NB, 0.87 and 1.00 for stage-level WR, 0.89 and 1.00 for stage-level WO when averaged across 2000 replicates, is non-negligible. Similar magnitudes of difference are also observed in the FS-AT test, indicating that both the FS and FS-AT tests are influenced by the correlation structure. We next discuss the underlying mechanism driving the observed decrease in power in simulation scenario S3.
@@
In simulation scenario S3, the marginal correlation between the survival time and time-to-hospitalization endpoints influences the chances to actually compare the time-to-hospitalization endpoint (conditioning on receiving ties in the survival time layer). As a result, although there is no treatment effect on the time-to-hospitalization endpoint marginally, the spurious, negative “treatment effects” are observed based on the pairwise comparisons between participants with no difference on the survival time layer. Such spurious negative “treatment effects” are the core reason for the reduced power in detecting the true positive treatment effects with both FS and FS-AT tests. These results underscore the importance of carefully considering the inclusion of additional endpoints in a hierarchical testing structure, as inter-endpoint correlations can affect treatment effect detection in unanticipated ways (verbeeck2019generalized). An endpoint that appears neutral with respect to power may, after conditioning on higher-layer outcomes, exhibit spurious negative or positive “treatment effects”, potentially affecting statistical inference. Additionally, Table 3 shows that the final stages of FS and FS-AT tests leave the same proportion of ties (3.15% in simulation scenario S3 and 1.34% in S4, respectively). This confirms that employing the last two stages of the FS-AT test prevents introducing additional ties and ensures full utilization of information, as in the standard FS test.
@@
Table 3: Stage-level decomposition of wins, losses, and ties in simulation scenarios S3 and S4 with 750-day follow-up time averaged across 2000 replicates. $\mathscr{D}_{ij}(d)$ and $\mathscr{T}_{ij1}(t_{1})$ represent comparison functions for survival time and time-to-hospitalization endpoints with thresholds $d$ and $t_{1}$, respectively. Similar for $\mathscr{D}_{ij}(0)$ and $\mathscr{T}_{ij1}(0)$. The stage-level NB/WO/WR with the bolded values correspond to those relevant to the time-to-hospitalization endpoint. Both simulation scenarios S3&S4 feature modest treatment effects on the survival time endpoint only, where simulation scenario S3 assumes a correlation between endpoints and simulation scenario S4 assumes no such correlation.
Scenario Method Stage Win(%) Tie(%) Loss(%) Stage-level
NB(%) WO WR
S3 FS $\mathscr{D}_{ij}(0)$ 37.29 35.15 27.56 9.72 1.22 1.35
$\mathscr{T}_{ij1}(0)$ 14.93 3.15 17.07 -6.10 0.89 0.87
FS-AT $\mathscr{D}_{ij}(d)$ 30.27 48.01 21.71 8.56 1.19 1.39
$\mathscr{T}_{ij1}(t_{1})$ 16.00 13.28 18.73 -5.69 0.89 0.85
$\mathscr{D}_{ij}(0)$ 2.48 8.76 2.04 3.29 1.07 1.21
$\mathscr{T}_{ij1}(0)$ 2.77 3.15 2.85 -0.90 0.98 0.97
S4 FS $\mathscr{D}_{ij}(0)$ 37.22 35.18 27.60 9.62 1.21 1.35
$\mathscr{T}_{ij1}(0)$ 16.90 1.34 16.94 -0.10 1.00 1.00
FS-AT $\mathscr{D}_{ij}(d)$ 30.20 48.04 21.76 8.43 1.18 1.39
$\mathscr{T}_{ij1}(t_{1})$ 18.31 11.39 18.34 -0.08 1.00 1.00
$\mathscr{D}_{ij}(0)$ 2.39 6.86 2.14 2.23 1.05 1.12
$\mathscr{T}_{ij1}(0)$ 2.76 1.34 2.76 -0.02 1.00 1.00
@@ 3.4 Influence of the caliper
In this subsection, we demonstrate the influence of caliper choices on adaptive thresholds. In addition to the default caliper ($c=20\%$), we consider $c=10\%,40\%$, and the combined calipers. The combined caliper is an eight-stage FS-MT test that progressively employs $c=40\%,20\%,10\%$ calipers: $\text{FS-MT}\left(q_{D}^{40\%},q_{1}^{40\%},q_{D}^{20\%},q_{1}^{20\%},q_{D}^{10\%},q_{1}^{10\%},0,0\right).$ The results are presented in Table 4. In this simulation, a larger caliper is more effective at bypassing random differences, while a smaller caliper is more sensitive to detecting treatment effects on the survival time layer. The combined caliper is slightly enhanced by the additional stages used to bypass random differences. When increasing the caliper from 10% to 20% and 40%, the performance gains in simulation scenarios S1, S2, S5, and S6, where treatment effects primarily occur on the time-to-hospitalization layer, generally outweigh the performance losses observed in the remaining scenarios where effects are concentrated on the survival time layer. This is expected since the adaptive thresholds work to keep detecting true signals while bypassing random differences. The combined caliper approach achieves the highest power under simulation scenarios S1, S2, S5, and S6, which is likely a result of its enhanced comparison structure that is more effective at bypassing the random differences. In general, the FS-AT test remains robust to the choice of caliper values as no rapid change in the performance of FS-AT test with different calipers is observed under most simulation scenarios.
Table 4: Empirical power (%) obtained by FS-AT with different calipers $c$ in simulation scenarios S1-S8. Notation: FU represents the schedule follow-up length; Caliper-10% refers to the FS-AT test with caliper $c=10\%$, similar for 20% and 40%; Combined Caliper refers to an eight-stage FS-AT test that progressively employs $c=40\%,20\%,10\%$ calipers.
@@ 3.5 Influence of the weight
In this subsection, we show the influence of weight choices on adaptive thresholds. Alternative to the default weight choice $w=1$, we also examine $w\in\{0.1,0.3,0.5\}$. In the resulting four candidate tests, FS-AT(1), FS-AT(0.5), FS-AT(0.3), FS-AT(0.1), a smaller $w$ places greater emphasis on the survival time endpoint. According to the results presented in Table 5, in simulation scenarios S1, S2, S5, and S6, where the treatment effects are primarily on the time-to-hospitalization layer, smaller $w$ leads to lower power. This is not surprising since emphasizing the higher-layer endpoint naturally reduces the chance in detecting the treatment effect on the lower layer. In simulation scenarios S3, S4, S7, and S8, where the treatment effects are primarily on the survival time layer, relatively more comparable power is achieved by different choices of $w$. These indicate that the FS-AT test is generally robust to the choice of $w$ in terms of maintaining the capacity to detect treatment effects on the survival time layer.
Table 5: Empirical power (%) obtained by FS-AT tests with different weights $w$ in simulation scenarios S1-S8. Notation: FU represents the schedule follow-up length; FS-AT(0.1) refers to the FS-AT test with weight $w=0.1$, similar for FS-AT(0.3), FS-AT(0.5), and FS-AT(1.0).
Scenario FU FS-AT(0.1) FS-AT(0.3) FS-AT(0.5) FS-AT(1.0)
@@ 4 Case Study: Digitalis Investigation Group (DIG) clinical trial
To illustrate the proposed testing procedure, we apply the FS-AT test to analyze the Digitalis Investigation Group (DIG) trial (ClinicalTrials.gov identifier: NCT00000476). This clinical trial investigated the effect of digoxin on patients with heart failure. Patients who had heart failure and a left ventricular ejection fraction of 0.45 or less were eligible for the primary randomization. This study was neutral for its primary outcome of all-cause mortality but evaluated both death and hospitalization events. When these two outcomes were analyzed separately, treatment effects appeared mainly on the hospitalization outcomes. Alternatively, we hereby employ FS-AT and standard FS tests to reanalyze this dataset. In this case study, both death and hospitalization outcomes are analyzed in the time-to-event form with the right censoring information taken into account (i.e., survival time and time to hospitalization). In developing the endpoint hierarchy, time to (all-cause) death is included as the higher layer, and time to (all-cause) hospitalization is included as the lower layer in both FS-AT and FS tests. For illustration of the differential operating characteristics between these tests, we focus on key subgroups from the trial population and use stratification for better adjustment in this analysis. That is, we study the subgroup population with New York Heart Association (NYHA) functional class III or IV and stratify by ejection fraction (<0.25<0.25 or 0.250.450.25-0.45), cause of heart failure (ischemic or nonischemic), and age (<70<70 or $\geq 70$). The number of patients within each stratum is presented in Supporting Information S4. For the FS-AT test, the adaptive thresholds with default weight and caliper ($w=1,c=20\%$) are adopted. In both approaches, the stratification step follows the details presented in finkelstein1999combining. Specifically, we first divide patients into distinct strata indexed by $r=1,2,\dots,R$ ($R=8$ in our case study with the DIG trial), and let $\mathcal{A}_{r}$ denote the set of indices for the $n_{r}$ participants in stratum $r$. For each individual $i\in\mathcal{A}_{r}$, the individual score is computed within stratum as $U_{i}=\sum_{j\in\mathcal{A}_{r}}U_{ij}.$ The stratified test statistic then takes the form $S_{\text{strat}}=\sum_{r=1}^{R}\sum_{i\in\mathcal{A}_{r}}Z_{i}U_{i},$ where $Z_{i}$ indicates treatment assignment. The variance of $S_{\text{strat}}$ is estimated analogously to the unstratified case by aggregating within-stratum variances: $\widehat{\text{Var}}(S_{\text{strat}})=\sum_{r=1}^{R}\frac{m_{r}(n_{r}-m_{r})}{n_{r}(n_{r}-1)}\left(\sum_{i\in\mathcal{A}_{r}}U_{i}^{2}\right),$ where $m_{r}$ is the number of treated participants in stratum $r$. A summary of alternative stratification approaches is provided in backer2025stratification.
@@
The resulting p-values are 0.033 and 0.055 for FS-AT and FS tests, respectively. Compared with the FS test, the FS-AT test yields a smaller p-value and concludes a significant treatment effect at the 0.05 level. To understand the underlying driver of such a difference, we present decomposition matrices and stage-level contribution to the overall hypothesis testing, quantified by NB, WO, and WR in Table 6. In this stratified analysis, the decomposition is obtained in a manner consistent with the stratified testing procedure. Specifically, within each stratum, we compute the counts of wins, ties, and losses from pairwise comparisons between treatment and control participants. These counts are then summed across all strata to yield the overall numbers of wins ($N_{+}$), ties ($N_{0}$), and losses ($N_{-}$). The corresponding proportions, expressed as percentages of the total number of pairwise comparisons between treatment and control groups, form the stage-level decomposition results presented in the table. These values are then used to calculate the stage-level GPC measures, which are reported to aid interpretation of each endpoint’s contribution to the overall test. It is noticeable that, the treatment effect is weak on the survival time endpoint but relatively stronger on the time-to-hospitalization endpoint. For instance, in the FS test, the stage-level NB is approximately -0.77% for the survival time endpoint and 5.09% for the time-to-hospitalization endpoint, indicating a noticeably stronger treatment effect in favor of the treatment group at the latter stage. A similar pattern is observed in the FS-AT test, where the corresponding NB values are -0.48% and 4.92% at the first two stages, when most ties were resolved. Consistent trends are also reflected in WO and WR. Therefore, the standard FS test may be less powerful as it assigns the strict priority given to the survival time. In contrast, the FS-AT test demonstrates a higher capacity in detecting the treatment effect under such cases. Finally, we report p-values for the hypothesis tests and interpret statistical significance at the 0.05 level only for the purposes of demonstration. Nevertheless, we recognize the caveat that, in practice, study conclusions may not depend solely on whether a p-value is greater than or less than a single cutoff point like 0.05.
Table 6: Stage-level decomposition of wins, losses, and ties in the DIG trial case study. $\mathscr{D}_{ij}(d)$ and $\mathscr{T}_{ij1}(t_{1})$ represent comparison functions for time to all-cause death and time to all-cause hospitalization endpoints with thresholds $d$ and $t_{1}$, respectively. Similar for $\mathscr{D}_{ij}(0)$ and $\mathscr{T}_{ij1}(0)$. The FS test has a p-value of 0.055 and the FS-AT test has a p-value of 0.033.
Method Stage Win(%) Tie(%) Loss(%) Stage-level
NB(%) WO WR
FS $\mathscr{D}_{ij}(0)$ 17.33 64.58 18.10 -0.77 0.98 0.96
$\mathscr{T}_{ij1}(0)$ 29.60 10.47 24.51 5.09 1.17 1.21
FS-AT $\mathscr{D}_{ij}(d)$ 11.20 77.13 11.67 -0.48 0.99 0.96
$\mathscr{T}_{ij1}(t_{1})$ 28.21 25.64 23.28 4.92 1.14 1.21
$\mathscr{D}_{ij}(0)$ 1.34 22.94 1.36 -0.02 1.00 0.99
$\mathscr{T}_{ij1}(0)$ 6.53 10.47 5.94 0.59 1.05 1.10
@@ 5 Discussion
In this study, we develop a hypothesis testing approach for composite time-to-event endpoints with a hierarchical structure. Modifying the standard FS test, the proposed FS-MT test adds layers with additional thresholds to pairwise comparisons. Weighted adaptive thresholds are also developed to reduce the potential challenge in determining the additional thresholds. With simulated datasets, we compare the performance of FS-AT and FS tests while addressing the influence of the follow-up time, treatment effect sizes, and correlations between the two endpoints. The numerical results show that the FS-AT test outperforms the FS test when the effect is primarily on the lower layer and generally maintains comparable power when the effect is primarily on the higher layer.
@@
In the FS-AT test, there are two tuning parameters: weight $w$ and caliper $c$. As a general rule of thumb, we recommend using $w=1,c=20\%$ as the default when clinical priors are not sufficiently strong. We have investigated several different choices of $c$ in the simulations to provide additional insights into their working mechanisms. In addition to different choices of $c$, the combined caliper approach is a strong alternative (e.g., $\text{FS-MT}\left(q_{D}^{40\%},q_{1}^{40\%},q_{D}^{20\%},q_{1}^{20\%},q_{D}^{10\%},q_{1}^{10\%},0,0\right)$ as employed in Section 3.4). The performance of the combined caliper approach tends to rely mainly on the largest caliper because the priority is given to this caliper by the specification of the hierarchy. It is also worth noting that extra attentions are given to the lower layer (i.e., time-to-hospitalization endpoint in our example) in this approach since each one of the added calipers will provide additional opportunities for the lower layer to contribute information to the pairwise comparison. In this combined caliper approach, the additional computation time brought by the added stages may need to be considered, particularly when the study sample size is large or repeated calculations are required (e.g., resampling-based inference). On the other hand, when $c$ is fixed, the weight $w\in[0,1]$ increases the adaptive threshold of the lower layer to address the higher layer endpoint. Since the reciprocal function in the $q_{1}^{c}/w$ structure is not on a linear scale, care is needed to use a relatively small $w$. Apart from adopting the reciprocal function, one may also consider the two-caliper structure ($c_{1},c_{2}$). This architecture employs a larger quantile to increase the lower layer threshold. When a small $w$ is desired, one may employ this approach and adjust $c_{2}$ to avoid the potential trivial stage when $w$ is specified close to zero. Despite the flexibility of the proposed adaptive threshold system, the structure ($w,c$) or ($c_{1},c_{2}$) should be pre-specified. Finally, when domain knowledge allows for precisely specifying all thresholds, directly implementing the FS-MT test could potentially enhance the clinical interpretability of each stage. In general, the FS-AT test serves to simplify the process for selecting tuning parameters based on general rule of thumbs, when clinical information is insufficient to determine the thresholds.
We note that, whether using the FS or FS-MT (FS-AT) test, clinical relevance remains important in selecting and prioritizing endpoints in the outcome specification stage. In general, the FS-MT test still assumes that composite endpoints and their hierarchy are defined a priori based on clinical judgment. That is, the FS-MT test does not completely overturn this ordering; rather, it introduces a more flexible comparison mechanism that starts from the highest-priority endpoint but then permits evaluation of lower-tier endpoints only when differences on the top layer are within a non-informative range. Hence, the FS-MT test occupies a middle ground between strict hierarchical enforcement (the standard FS test) and complete re-ranking of endpoints. In practice, if strong clinical priors exist for fixed thresholds, they should be incorporated within the FS-MT test. In the absence of such priors, the FS-AT test offers a pragmatic alternative to adaptively define these thresholds. We have not argued the use of FS-MT and FS-AT tests to replace the standard FS test in all settings. In fact, when clinical relevance demands strict hierarchical testing, the standard FS test will still be preferred as primary analysis, but FS-MT and FS-AT tests may contribute useful information in secondary analysis. In any case, large discrepancies between standard FS and FS-MT (FS-AT) testing results should prompt reflection on what aspect of treatment benefit is most meaningful given the context of the specific clinical trial.
Although our focus has been on hypothesis testing with the FS test, the proposed comparison structure can also be used to estimate treatment effect measures within the GPC framework. By incorporating the developed win functions into the computation of $_{+}, NN_{-}$, and $N_{0}$, one can obtain corresponding treatment effect estimates, as shown in equation (1). This formulation naturally extends beyond the NB, WO, and WR measures and can be adapted to other measures and methods within the GPC framework. While our study focuses on the FS test, we acknowledge that alternative hypothesis testing procedures within the GPC framework possess distinct operating characteristics and may outperform each other under certain conditions. For instance, to better handle censoring in time-to-event endpoints, the proposed win function could be incorporated into alternative approaches such as inverse probability of censoring weighting (dong2020inverse; dong2021adjusting) or other methods specifically developed for censored data (peron2018extension; deltuvaite2023generalized; deltuvaite-thomas2025rightcensored). Addressing censoring in our setting with multiple thresholds is a valuable topic for future research.